# Quantum simulation of one-dimensional fermionic systems with Ising Hamiltonians

### v1

@@ Abstract
In recent years, analog quantum simulators have reached unprecedented quality, both in qubit numbers and coherence times. Most of these simulators natively implement Ising-type Hamiltonians, which limits the class of models that can be simulated efficiently. We propose a method to overcome this limitation and simulate the time-evolution of a large class of spinless fermionic systems in 1D using simple Ising-type Hamiltonians with local transverse fields. The time complexity of the simulation scales with the square root of the inverse error, and thus favorably compared to the worst-case error of first-order product formulas. Our method is based on domain wall encoding [1], which is implemented via strong (anti-)ferromagnetic couplings $|J|$. We show that in the limit of strong $|J|$, the domain walls behave like spinless fermions in 1D. The Ising Hamiltonians are one-dimensional chains with nearest-neighbor and, optionally, next-nearest-neighbor interactions. As a proof-of-concept, we perform numerical simulations of various 1D-fermionic systems using domain wall evolution and accurately reproduce the systems’ properties, such as topological edge states, Anderson localization, quantum chaotic time evolution and time-reversal symmetry breaking via Floquet-engineering. Our approach makes the simulation of a large class of fermionic many-body systems feasible on analogue quantum hardware that natively implements Ising-type Hamiltonians with transverse fields.
@@ I Introduction
The work is structured as follows: first, we will introduce the theory of how domain wall encoding [1] can be used to simulate spinless fermionic systems in one dimension to an arbitrarily close approximation. Then, we will numerically investigate the static properties of the approximation on the example of the SSH model [31], a key model for understanding topological excitations [32]. To investigate the simulation of dynamic properties, we will discuss simulations of the Aubry-Andre model [33], which has an extended-localized phase transition similar to the Anderson model. Furthermore, we investigate the influence of decoherence on the results by reproducing the dynamical phase transitions of the AA-model. We test the method’s ability to simulate interactions between particles in the inhomogeneous Heisenberg XXZ model and, lastly, we make use of Floquet engineering to realize exotic imaginary next-nearest-neighbor hopping terms.
@@ II.2 Domain walls as interacting particles
Additionally, the excitation conservation argument also holds for one-dimensional spin chains with periodic boundary conditions, if the Ising chain is also periodic. In this case the number of particles is forced to be even, since on a periodic chain, the number of domain walls has to be even. This limitation, however, can be overcome by doubling the Ising chain by appending a copy of itself, before closing the periodic boundary. To simulate time evolution, the initial state has to be prepared in each subsystem respectively. This can be thought of as explicitly simulating two periods of the periodic model, instead of simulating one period explicitly and the infinite copies implicitly. The doubling of the particle number allow to simulate odd particles numbers in the original system with periodic boundaries. In case of periodic boundary conditions, some care has to be taken when preparing the initial state due to symmetry considerations. We provide more details on this question in appendix A.
@@
While in this work we assume ferromagnetic couplings $J<0$, note that our proposal also works for devices where strong anti-ferromagnetic couplings are available exclusively, or are favorable. If one exchanges the logical meaning of 0 and 1 on every other qubit and at the same time changes the sign of the nearest-neighbor couplers, the Hamiltonian Eq. (17) can be realized with $J>0$.
Figure 1: Illustration of the energy landscape of Hamiltonian $H_{DW}$ in Eq. (10) and Eq. (17). As the number of domain walls increases, the energy landscape moves upwards by $4|J|$. For sufficiently large $|J|$, the subspaces corresponding to different numbers of domain walls are separated by energy gaps of $\approx 4|J|$ (red shaded areas). The transitions across these energy gaps are neglectable when considering time-evolution according to $H_{DW}$, however, the subspaces seem to perturb each other and alter the time-evolution within each subspace. At a linear time-overhead, this perturbation can be made arbitrarily small.
@@ III.1 Perturbative analysis via Schrieffer-Wolff transformation
where $V$ is the block-off-diagonal perturbation and $D$ a diagonal matrix that collects the energy offsets. $H_{\text{Fermi}}^{\text{odd}}$ denotes the fermionic Hamiltonian restricted to odd particle numbers. Note that here we discuss the setting of an odd number of domain walls, but the analysis extends readily to the setting of an even number of particles. Since the energy shift $D$ is constant in each block, we can ignore $D$ in the analysis of the dynamics within the respective subspaces. However, the energy shifts allow us to argue that the effect of the perturbation $V$ vanishes for large $|J|$. We can diagonalize the
@@
where $\Lambda$ is the diagonal matrix of eigenvalues and $U$ the unitary of eigenvectors. $U$ is also block-diagonal and we denote the projections into the respective subspaces as $U_{M}$. In this basis, the perturbations transform into $\tilde{V}=U^{\dagger}VU$, which remain block-off-diagonal where the entries are given by
@@
Note that the eigenbasis $U_{M}$ of $H_{DW}^{(M)}$ also diagonalizes $H_{\text{Fermi}}^{(M)}$ and the eigenvalues are the same up to the energy shift $(N-2M)J-\frac{1}{4}\sum_{n=1}^{N-1}v_{n}$.
@@ III.2 Parameter re-scaling and time overhead
Compare this to the error bound of the first-order product formula [38], also known as Trotterization [15], used to simulate continuous time dynamics on gate-based quantum simulators. The gate complexity scales with $O(T^{2}/\varepsilon)$. We find that our approach of simulating continuous time evolution with an analog quantum device displays the same asymptotic behavior with respect to the time $T$. However, our quantum simulation algorithm scales only with the square root of the inverse error $\varepsilon$. Thus, our approach improves upon the first-order Trotterization error bound.
@@ IV Static analysis: the SSH-model
We will investigate the static properties of the effective Hamiltonians $H_{DW}^{(M)}$ using the example of the Su-Schrieffer-Heeger (SSH) model [31, 32]. The SSH model is a one-dimensional model on a linear chain, where the unit cell contains two sites $A$ and $B$. A particle can hop from a site $A$ to site $B$ in either the same or in the previous unit cell. The key feature of the SSH model are the modulated hopping energies within the same unit cell and between two distinct unit cells, denoted by $v$ and $w$ respectively.
@@
The SSH model is one of the simplest model displaying a topological phase transition. For $|v|>|w|$, the system is in a topological trivial phase with a band gap, while for $|v|<|w|$ the system is topologically non-trivial. In the topologically non-trivial phase, there are zero-energy edge states sitting in the band gap. These edge states are localized on the ends of the SSH chain. The two topological phases are separated by a conducting phase $|v|=|w|$. The SSH model has been simulated on quantum simulators based on atom-optics [32], superconducting transmon qubits [41] and quantum dots [42].
@@ V.1 Single particle
First we compare the time evolution induced by the Hamiltonian $H_{\text{Fermi}}$ and $H_{DW}$ in the respective subspace. We simulate a system of $N=13$ sites, initialize the particle on site 7 and simulate unitary time evolution according to the Hamiltonians $H_{\text{Fermi}}$ and $H_{DW}$. In Figure 3, we plot the occupations of each site
@@
where $\langle O\rangle_{T}$ denotes the expectation value $\langle\psi_{T}|O|\psi_{T}\rangle$. The site populations defined in Eq. (52) are also valid in the multi-particle setting. Note that the required expectation values are at most two-local and are thus easily accessible via measurements in the computational basis, e.g. measurement of the persistent current in superconducting flux qubits. By comparing Figure LABEL:fig:TimeExact and LABEL:fig:TimeApprox, we can see that the time-evolution of the site occupations is accurately captured by the domain wall approximation in both phases of the model, as well as at the critical point.
@@
Figure 4: Subspace fidelity $\mathcal{F}_{\text{subspace}}$ (top) and state fidelity $\mathcal{F}$ (bottom) for different values of the coupling $J$. As for the simulation shown in Figure 3, the system is initialized with a single particle on site 7 and evolves under an Aubry-Andre Hamiltonian in the localized, critical and extended phase respectively.
@@
We can conclude that even at reasonable coupling strengths of $|J|/|t|\approx 5$, the domain wall approximation closely reproduces the dynamics of the Aubry-Andre model for a single particle. The extended simulation in Figure LABEL:fig:TimeApprox suggest that information can spread several time through the entire system on a time scale with reasonable fidelities and that many interesting properties of the original system’s dynamics can still be probed using the domain wall simulator. Next, we investigate the dynamics of multiple domain walls.
@@ V.2 Multiple spinless fermions
Here we will not only consider different values for the on-site disorder $\lambda$ in Eq. (49), but also introduce disorder in the hopping energies $t_{n}$ via the parameter $\mu$ in Eq. (50).
@@ V.3 Dynamical quantum phase transition
Usually, DQPTs are observed if the quench dynamics is across a quantum phase transition. We gauge the fidelity by numerically solving the Schrödinger equation both for the the Aubry-Andre model for spins and domain walls. Additionally, we investigate the impact of decoherence on the quantum simulation by solving the Lindblad master equation
@@
Figure 7: Rate function $r_{\text{exact}}$ for the exact time evolution, the domain wall approximation $r_{DW}$ and the domain wall approximation with noise $r_{DW}^{*}$; the initial state in Fock space are $|00000100000\rangle$ ($M=1$), $|01000100010\rangle$ ($M=3$) or $|01010101010\rangle$ ($M=5$), while the Hamiltonian is the Aubry-Andre model in the critical phase with $N=11$, $t=1$, $\lambda=2$, $\mu=1.5$; the noise is modeled by relaxation with $\Gamma_{1}=1/500$ and dephasing with $\Gamma_{2}=1/100$ of each individual qubit in the simulator. The results shown here were obtained using the re-scaling Eq. (43).
@@
This raises the question how feasible the simulation is on real devices. As stated, this will depend on the concrete implementation, but here we will give some estimates for realistic system parameters. If we identify a unit of energy in our simulation with GHz, then a unit of time corresponds to ns. Hence the experiment discussed in this section requires qubits with relaxation time $T_{1}=500$ ns and dephasing time $T_{2}=100$ ns, which are conservative estimates for contemporary superconducting flux qubits [49]. The ferromagnetic coupling energy in the simulations corresponds to 5 GHz, which is also state of the art [17]. At the same time, with $\mu=1.5$ and hopping energies $t_{n}$ from Eq. (50), the energy splitting of the individual qubits becomes very small, such that thermal noise starts to increase the thermal decay rates [50]. We leave it to further research how this issue can be mitigated, but we expect that the critical phenomena discussed here can be probed with contemporary quantum hardware.
@@ VI Interacting system: inhomogeneous XXZ model
The models we have analyzed so far only required hopping- and on-site energies, but no interactions between neighboring particles. In order to validate the implementation of the particle interactions, we simulate the inhomogeneous Heisenberg XXZ model.
@@
In line with the high operator fidelities in Figure 10 (a), the dynamical properties are well reproduced. The survival probability $\mathcal{P}$ of the domain wall approximation coincides closely with the solution of the XXZ model, as seen in Figure 10 (b), while the fidelities in Figure 10 (d) remain $>0.5$ throughout the simulation. The fidelity of the simulation increases with increasing $\theta$. This is a consequence of the energy re-scaling to effectively increase the domain wall coupling, such that the $M$-subspaces of $H_{DW}$ remain well separated even for large $\theta$. Note, that we use the standard re-scaling Eq. (43) for all models here and that the fidelities can be improved arbitrarily at a linear time-overhead.
@@ VII Time-dependent Hamiltonian: Next-nearest-neighbor hopping via Floquet engineering
The time evolution for a integer multiple of the period $T_{\text{evol}}=m\tau$ is described by
@@
Due to the truncation of the Magnus expansion, the Floquet Hamiltonian only describes the time evolution approximately and for a evolution time $T_{\text{evol}}$ the error is [54]
$\varepsilon_{\text{Floquet}}=O(\tau^{2}T_{\text{evol}})\ .$ (84)
@@
The re-scaling also applies for time-dependent parameters. However, if we maintain the original frequency $\Omega$, we effectively simulate the same time evolution, but with a larger frequency $\alpha\Omega$ and the Floquet error is reduced for $\alpha>1$
$\varepsilon_{\text{Floquet}}=O(\tau^{2}\alpha^{-2}T_{\text{evol}})\ .$ (86)
@@
Thus, we can make the simulation accurate up to an error $\varepsilon$ by choosing
$\alpha=O\left(\frac{1}{\sqrt{\varepsilon}}\max\left\{\tau\sqrt{T_{\text{evol}}},\frac{T_{\text{evol}}}{\tau|J|}\right\}\right)\ .$ (89)
Note that this re-scaling is useful for quantum devices where the drive frequency $\Omega=\tau^{-1}$, the coupling $|J|$ or both cannot be increased arbitrarily. In summary, the evolution under $H_{NNN}$ for time $T_{\text{evol}}$ can be simulated to arbitrary precision $\varepsilon$ using the Ising-type Hamiltonian $H_{DW}$ with modulated transverse fields in time at most $O(T_{\text{evol}}^{2}/\sqrt{\varepsilon})$.
@@
We use a strong ferromagnetic domain wall coupling of $J=-15$. This is necessary for this approach, since now the energy scale of the fermionic system is not only determined by the hopping energies $|t_{n}^{(0)}|+|t_{n}^{(1)}|$, but also by the drive frequency $\Omega$. Therefore, a strong domain wall coupling is required to keep the $M$-subspaces well separated. Alternatively, one could re-scale the system parameters and the simulation time to achieve the same accuracy, as discussed previously.
[…]

### v2

@@ Abstract
In recent years, analog quantum simulators have reached unprecedented quality, both in qubit numbers and coherence times. Most of these simulators natively implement Ising-type Hamiltonians, which limits the class of models that can be simulated efficiently. We propose a method to overcome this limitation and simulate the time-evolution of a large class of spinless fermionic systems in 1D using simple Ising-type Hamiltonians with local transverse fields. Our method is based on domain wall encoding [1], which is implemented via strong (anti-)ferromagnetic couplings $|J|$. We show that in the limit of strong $|J|$, the domain walls behave like spinless fermions in 1D. The Ising Hamiltonians are one-dimensional chains with nearest-neighbor and, optionally, next-nearest-neighbor interactions. As a proof-of-concept, we perform numerical simulations of various 1D-fermionic systems using domain wall evolution and accurately reproduce the systems’ properties, such as topological edge states, Anderson localization, quantum chaotic time evolution and time-reversal symmetry breaking via Floquet-engineering. Our approach makes the simulation of a large class of fermionic many-body systems feasible on analogue quantum hardware that natively implements Ising-type Hamiltonians with transverse fields.
@@ I Introduction
The work is structured as follows: first, we will introduce the theory of how domain wall encoding [1] can be used to simulate spinless fermionic systems in one dimension to an arbitrarily close approximation. Then, we will numerically investigate the static properties of the approximation on the example of the Su-Schrieffer-Heeger (SSH) model [31], a key model for understanding topological excitations [32]. To investigate the simulation of dynamic properties, we will discuss simulations of the Aubry-Andre model [33], which has an extended-localized phase transition similar to the Anderson model. Furthermore, we investigate the influence of decoherence on the results by reproducing the dynamical phase transitions of the Aubry-Andre model. We test the method’s ability to simulate interactions between particles in the inhomogeneous Heisenberg XXZ model and, lastly, we make use of Floquet engineering to realize exotic imaginary next-nearest-neighbor hopping terms.
@@ II.2 Domain walls as interacting particles
Additionally, the excitation conservation argument also holds for one-dimensional spin chains with periodic boundary conditions, if the Ising chain is also periodic. In this case the number of particles is forced to be even, since on a periodic chain, the number of domain walls has to be even. This limitation, however, can be overcome by doubling the Ising chain by appending a copy of itself, before closing the periodic boundary. To simulate time evolution, the initial state has to be prepared in each subsystem respectively. This can be thought of as explicitly simulating two periods of the periodic model, instead of simulating one period explicitly and the infinite copies implicitly. The doubling of the particle number allows to simulate odd particle numbers in the original system with periodic boundaries. In case of periodic boundary conditions, some care has to be taken when preparing the initial state due to symmetry considerations. We provide more details on this question in appendix A.
@@
While in this work we assume ferromagnetic couplings $J<0$, note that our proposal also works for devices where strong anti-ferromagnetic couplings are available exclusively, or are favorable. If one exchanges the logical meaning of 0 and 1 on every other qubit and at the same time changes the sign of the nearest-neighbor couplers, the Hamiltonian Eq. (17) can be realized with $J>0$. In case of periodic boundary conditions, this requires the number of qubits to be even, which, again, can be ensured by doubling the chain.
Figure 1: Illustration of the energy landscape of Hamiltonian $H_{DW}$ in Eq. (10) and Eq. (17). As the number of domain walls increases, the energy landscape moves upwards by $4|J|$. For sufficiently large $|J|$, the subspaces corresponding to different numbers of domain walls are separated by energy gaps of $\approx 4|J|$ (red shaded areas). The transitions across these energy gaps are neglectable when considering time-evolution according to $H_{DW}$, however, the subspaces perturb each other and alter the time-evolution within each subspace. At a linear time-overhead, this perturbation can be made arbitrarily small.
@@ III.1 Perturbative analysis via Schrieffer-Wolff transformation
where $V$ is the block-off-diagonal perturbation and $D$ a diagonal matrix that collects the energy offsets. $H_{\text{Fermi}}^{\text{odd}}$ denotes the fermionic Hamiltonian restricted to odd particle numbers. Note that here we discuss the setting of an odd number of domain walls, but the analysis extends readily to the setting of an even number of particles. Since the energy shift $D$ is constant in each block, we can ignore $D$ in the analysis of the dynamics within the respective subspaces. However, the energy shifts allow us to argue that the effect of the perturbation $V$ vanishes for large $|J|$. We can diagonalize the block-diagonal
@@
where $\Lambda$ is the diagonal matrix of eigenvalues and $U$ the unitary of eigenvectors. $U$ is also block-diagonal and we denote the projections into the respective subspaces as $U_{M}$. In this basis, the perturbation transforms into $\tilde{V}=U^{\dagger}VU$, which remains block-off-diagonal where the entries are given by
@@
Note that the eigenbasis $U_{M}$ of $H_{\text{Fermi}}^{(M)}$ also diagonalizes $H_{DW}^{(M)}$ and the eigenvalues are the same up to the energy shift $(N-2M)J-\frac{1}{4}\sum_{n=1}^{N-1}v_{n}$.
@@
Note that in the literature it is common to use the spectral norm of two time evolution operators as a metric for the error of a quantum simulation. Here, we consider the infidelity $\varepsilon=1-\mathcal{F}$. Unlike the spectral norm, the infidelity is insensitive to global phase factors. Since our method inevitably produces global phases due to the energy offsets, the infidelity is the more natural metric to gauge its accuracy. However, the spectral norm could also be considered after correcting for the known global phases and ultimately both errors are equivalent.
@@ III.2 Parameter re-scaling and time overhead
Compare this to the error bound of the first-order product formula [38], also known as Trotterization [15], used to simulate continuous time dynamics on gate-based quantum simulators. The gate complexity scales with $O(T^{2}/\varepsilon_{\|\cdot\|})$, where $\varepsilon_{\|\cdot\|}$ indicates that the error in Ref.[38] is computed in terms of the spectral norm. However, $\varepsilon=O(\varepsilon_{\|\cdot\|}^{2})$, as discussed in appendix C. Thus, we find that our approach of simulating continuous time evolution with an analog quantum device displays the same asymptotic behavior with respect to time $T$ and error $\varepsilon$.
@@ IV Static analysis: the SSH-model
We will investigate the static properties of the effective Hamiltonians $H_{DW}^{(M)}$ using the example of the SSH model [31, 32]. The SSH model is a one-dimensional model on a linear chain, where the unit cell contains two sites $A$ and $B$. A particle can hop from a site $A$ to site $B$ in either the same or in the previous unit cell. The key feature of the SSH model are the modulated hopping energies within the same unit cell and between two distinct unit cells, denoted by $v$ and $w$ respectively.
@@
The SSH model is one of the simplest models displaying a topological phase transition. For $|v|>|w|$, the system is in a topological trivial phase with a band gap, while for $|v|<|w|$ the system is topologically non-trivial. In the topologically non-trivial phase, there are zero-energy edge states located in the band gap. These edge states are localized on the ends of the SSH chain. The two topological phases are separated by a conducting phase $|v|=|w|$. The SSH model has been simulated on quantum simulators based on atom-optics [32], superconducting transmon qubits [41] and quantum dots [42].
@@ V.1 Single particle
First we compare the time evolution induced by the Hamiltonian $H_{\text{Fermi}}$ and $H_{DW}$ in the respective subspace. We simulate a system of $N=13$ sites, initialize the particle on site 7 and simulate unitary time evolution according to the Hamiltonians $H_{\text{Fermi}}$ and $H_{DW}$. We do not apply the standard re-scaling Eq. (43) in order to obtain an intuition of the impact of the domain wall coupling $|J|$. In Figure 3, we plot the occupations of each site
@@
where $\langle O\rangle_{T}$ denotes the expectation value $\langle\psi_{T}|O|\psi_{T}\rangle$. The site populations defined in Eq. (52) are also valid in the multi-particle setting. Note that the required expectation values are at most two-local and are thus easily accessible via measurements in the computational basis, e.g. measurement of the persistent current in superconducting flux qubits. By comparing Figure 3 top and bottom row, we can see that the time-evolution of the site occupations is accurately captured by the domain wall approximation in both phases of the model, as well as at the critical point.
@@
Figure 4: Subspace fidelity $\mathcal{F}_{\text{subspace}}$ (top) and state fidelity $\mathcal{F}$ (bottom) for different values of the coupling $J$. As for the simulation shown in Figure 3, the system is initialized with a single particle on site 7 and evolves under an Aubry-Andre Hamiltonian in the localized, critical and extended phase respectively. As in Figure 3, no re-scaling of the model parameters and time is applied and the results shown are for the real evolution time.
@@
We can conclude that even at reasonable coupling strengths of $|J|/|t|\approx 5$, the domain wall approximation closely reproduces the dynamics of the Aubry-Andre model for a single particle. The extended simulation in Figure 3 bottom row suggest that information can spread several time through the entire system on a time scale with reasonable fidelities and that many interesting properties of the original system’s dynamics can still be probed using the domain wall simulator. Next, we investigate the dynamics of multiple domain walls.
@@ V.2 Multiple spinless fermions
Here we will not only consider different values for the on-site disorder $\lambda$ in Eq. (49), but also introduce disorder in the hopping energies $t_{n}$ via the parameter $\mu$ in Eq. (50). Note that in the following simulations we employ the standard rescaling Eq. (43).
@@ V.3 Dynamical quantum phase transition
Usually, DQPTs are observed if the quench dynamics is across a quantum phase transition. We gauge the fidelity by numerically solving the Schrödinger equation both for the the Aubry-Andre model for spinless fermions and domain walls. Additionally, we investigate the impact of decoherence on the quantum simulation by solving the Lindblad master equation
@@
Figure 7: Rate function $r_{\text{exact}}$ for the exact time evolution, the domain wall approximation $r_{DW}$ and the domain wall approximation with noise $r_{DW}^{*}$; the initial state in Fock space are $|00000100000\rangle$ ($M=1$), $|01000100010\rangle$ ($M=3$) or $|01010101010\rangle$ ($M=5$), while the Hamiltonian is the Aubry-Andre model in the critical phase with $N=11$, $\lambda=2$, $\mu=1.5$; the noise is modeled by relaxation with $\Gamma_{1}=1/500$ and dephasing with $\Gamma_{2}=1/100$ of each individual qubit in the simulator. The results shown here were obtained using the re-scaling Eq. (43).
@@
This raises the question how feasible the simulation is on real devices. As stated, this will depend on the concrete implementation, but here we will give some estimates for realistic system parameters. If we identify a unit of energy in our simulation with GHz, then a unit of time corresponds to ns. Hence the experiment discussed in this section requires qubits with relaxation time $T_{1}=500$ ns and dephasing time $T_{2}=100$ ns, which are conservative estimates for contemporary superconducting flux qubits [49]. The ferromagnetic coupling energy in the simulations corresponds to 5 GHz, which is also state of the art [17]. At the same time, with $\mu=1.5$ and hopping energies $t_{n}$ from Eq. (50), the energy splitting of the individual qubits becomes small, if the re-scaling Eq. (43) is applied. Consequently, thermal noise starts to increase with the thermal decay rates [50]. While this may not be a severe concern for short simulations, we leave it to further research how this issue can be mitigated. However, we expect that the critical phenomena discussed here can be probed with contemporary quantum hardware.
@@ VI Interacting system: inhomogeneous XXZ model
The models we have analyzed so far only required hopping and on-site energies, but no interactions between neighboring particles. In order to validate the implementation of the particle interactions, we simulate the inhomogeneous Heisenberg XXZ model.
@@
In line with the high operator fidelities in Figure 10 (a), the dynamical properties are well reproduced. The survival probability $\mathcal{P}$ of the domain wall approximation coincides closely with the solution of the XXZ model, as seen in Figure 10 (b), while the fidelities in Figure 10 (d) remain $>0.5$ throughout the simulation. The fidelity of the simulation increases with $\theta$. This is a consequence of the standard energy re-scaling Eq. (43) to effectively increase the domain wall coupling, such that the $M$-subspaces of $H_{DW}$ remain well separated even for large $\theta$. Note, that by using the standard re-scaling Eq. (43) the fidelities can be improved arbitrarily at a linear time-overhead.
@@ VII Time-dependent Hamiltonian: Next-nearest-neighbor hopping via Floquet engineering
The time evolution for an integer multiple of the period $T_{\text{evol}}=m\tau$ is described by
@@
Due to the truncation of the Magnus expansion, the Floquet Hamiltonian only describes the time evolution approximately and for a evolution time $T_{\text{evol}}$ the error in terms of the infidelity is
$\varepsilon_{\text{Floquet}}\leq O(\tau^{4}T_{\text{evol}}^{2})\ .$ (84)
For the expression of the infidelity in Eq. (84) we use the error in terms of the spectral norm from Ref. [54], Eq. (22), and adapt it in accordance with the discussion in appendix C.
@@
The re-scaling also applies for time-dependent parameters. However, if we maintain the original frequency $\Omega$, we effectively simulate the same time evolution, but with a larger frequency $\alpha\Omega$ and the truncation error of the Magnus expansion is reduced. However, at the same time, we also scale the factor in the first equality of Eq. (81). Combining these contributions, the infidelity of the simulation for $\alpha>1$ is
$\varepsilon_{\text{Floquet}}=O(\tau^{4}\alpha^{-2}T_{\text{evol}}^{2})\ .$ (86)
@@
Thus, by considering Eq. (82), Eq. (85) and Eq. (86), we can make the simulation accurate up to an error $\varepsilon$ by choosing
$\alpha=O\left(\max\left\{\frac{\tau^{2}T_{\text{evol}}}{\sqrt{\varepsilon}},\frac{T_{\text{evol}}}{\tau|J|\sqrt{\varepsilon}}\right\}\right)\ .$ (89)
From Eq. (89) it is clear that the dominant error depends on the frequency and the domain wall coupling $|J|$. Note that this re-scaling is useful for quantum devices where the drive frequency $\Omega=\tau^{-1}$, the coupling $|J|$ or both cannot be increased arbitrarily. In summary, the evolution under $H_{NNN}$ for time $T_{\text{evol}}$ can be simulated to arbitrary precision $\varepsilon$ using the Ising-type Hamiltonian $H_{DW}$ with modulated transverse fields in time at most $O(T_{\text{evol}}^{2}/\sqrt{\varepsilon})$.
@@
We use a strong ferromagnetic domain wall coupling of $J=-15$. This is necessary for this approach, since now the energy scale of the fermionic system is not only determined by the hopping energies $|t_{n}^{(0)}|+|t_{n}^{(1)}|$, but also by the drive frequency $\Omega$, which could excite the system. Therefore, a strong domain wall coupling is required to keep the $M$-subspaces well separated. Alternatively, one could re-scale the system parameters and the simulation time to achieve the same accuracy, as discussed previously.
We extend the numerical analysis by performing the same simulation, but with initial states with multiple particles. We compute the fidelity of the simulation as a function of the stroboscopic evolution time $T_{\text{evol}}=m\tau$ with $m\in\mathbb{N}$, when the time evolution of the driven system coincides with the time evolution under the effective Hamiltonian. The result is shown in Figure 13 (b). We find that while the fidelity $\mathcal{F}$ of the simulation for $M=1$ is remarkably high throughout the simulation, for $M>1$, $\mathcal{F}$ decays substantially faster, but remains reasonably large over the simulated time interval, reaching $\approx 0.74$ at worst.